TEX and TUG NEWS

Mission Statement

The TEX Users Group (TUG) provides leadership:

1. to encourage and expand the use of TEX, , and related systems

2. to ensure the integrity and portability of TEX, , and related systems

3. to foster innovation in high-quality electronic document preparation

TEX and TUG NEWS is a newsletter for TEX and L^{A}TEX users alike: a forum for exchanging information, tips and suggestions; a regular means of communicating news items to one another; a place where information about TEX and TUG can be quickly disseminated.

Throughout the newsletter “TEX” is understood to mean TEX, L^{A}TEX, AMS-TEX, and other related programs and macros. TEX and TUG NEWS is produced with the standard L^{A}TEX distribution, and is to be as portable a document as possible.

The entire contents of this newsletter are being placed in the public domain. The source file of this issue will be placed in the aston, shsu, and stuttgart archives, as well as at the heidelberg, labrea, and ymir archives. Copying and reprinting are encouraged; however, an acknowledgement specifying TEX and TUG NEWS as the source would be appreciated.

Submissions to TEX and TUG NEWS should be short, the macros must work, and the files must run without special font or graphics requirements: this is to be a portable newsletter (the new font naming scheme has not yet been implemented). Correspondence may be sent via e-mail to tug@math.ams.com with the subject line NEWSLETTER. Regular mail should be addressed to the Publications Committee, c/o TUG Office at the address below.

TEX Users Group  Phone: 401-751-7760
P.O. Box 9506  FAX: 401-751-1071
Providence, RI 02940-9506 USA E-mail: tug@math.ams.com

TEX is a trademark of the American Mathematical Society

Editorial

A financial deficit for TUG, which became quite severe in the second half of 1991, forced the cancellation of TEX and TUG NEWS till now; as well, the July Proceedings were spread across two issues of TUGboat, in order to cut production and editorial costs. You will soon be receiving the first 1992 issue of TUGboat: 13, no. 1. We all hope that this financial crisis is now behind us. But you can help by checking that your institution has renewed its membership … And remember that the membership campaign asks you to TUG a member!

So, here we are, back on the air. And after such a long delay, there’s a lot of news to catch up on. But first: 23 March 1992 is the deadline for submission of entries for the 1992 Knuth Scholarship. See p. 3 of the newsletter for information, as well as pp. 563–64 of TUGboat 12, no. 4.

Other important TUG events have been the elections of last fall (p. 2), and some long-awaited information on the L^{A}TEX3 Project (pp. 4–6). And look at the opposite page — on the inside front cover, we now carry the new TEX Users Group Mission Statement. This is an important set of statements, which we will use to guide our direction, and our decisions, as we move through the 90s, and see the developments in (L^{A})TEX and their application to the production of fine typesetting.

We have a feature report on the state of TEX in Mexico (pp. 7–11), where a large meeting of the new Mexican TEX user group will convene towards the end of February. There are reports from other user groups (two new ones!), and meetings that have taken place and some to take place. There are two new columns — Typographer’s Inn with Peter Flynn, and New Publications, which is looking for a regular columnist. We hope to see new publications listed here in TEX and TUG NEWS, and reviews to appear in TUGboat, where there’s more room!

So, go read about what’s been happening — and send in your news, your views, and what works!

* * *

It is with sadness, both personal and professional, that we must also mark the passing of two very committed TUG members last year: Cathy Booth, who was intensely involved with the TEX users in Britain, and Sam Whidden, a founding member of the TEX Users Group. I was only one of many, many others who knew them, and who will not forget them.

Christina Thiele

Editor, TEX and TUG NEWS

First TUG Election by Mail

During 1991, a process was set in motion that would culminate in the first general election, by mail, for the TUG Board of Directors.

Twenty seven candidates were on the ballot to fill 15 positions on the Board. There were 726 ballots cast, of which 9 were disqualified. The newly elected Board members are:

Barbara Beeton  551 Anita Hoover  289 David Kellerman 345
Luzia Dietsche  377 Mimi Jett  284 Nico Poppelier  342
Kenneth Dreyhaupt 279 Y. Haralambous 380 Jon Radel  269
Michael Ferguson  299 Doug Henderson 357 Christina Thiele 354
Peter Flynn  328 Alan Hoenig  410 William Woolf  313

Of this new Board, five (Beeton, Henderson, Hoenig, Kellerman, and Thiele) have served on the Board previously. In addition to this list, the International Vice Presidents (Peter Abbott, the new Chair of the UK TUG, has replaced Malcolm Clark; Bernard Gaulle, GUTenberg; Roswitha Graham, Nordic TUG; Kees van der Laan, NTG; and Joachim Lammarsch, DANTE) remain on the Board. Malcolm Clark will be serving as TUG President for 1992.

Three honorary members also remain associated with the Board: Donald Knuth, Hermann Zapf, and Ray Goucher.

Unfortunately, we do not have any demographic information in time for this issue of TEX and TUG NEWS. We will pursue this information and report as soon as it becomes available.

There were some problems, as you might expect, with this first election. Probably the most notable was not enough time allowed for mail delivery outside of North America. This will be corrected in the next election.

I hope you will join me in thanking the following outgoing Board members for their hard work at establishing TUG, helping TUG and TEX to grow and become a success. (The number of years on the Board is shown in parentheses following each name.)

Lance Carnes^{2} (10) David Fuchs  (11) Patrick Ion  (7)
Bart Childs^{1,2}  (8) Regina Girouard (3) David Kratzer  (2)
John Crawford^{2} (7) Dean Guenther^{2} (5) Pierre MacKay^{1,2} (8)
Allen Dyer  (4) Hope Hamilton  (2) Craig Platt^{2}  (6)

Notes: ^{1}Former TUG President; ^{2}Present or former site coordinator.

Dean Guenther, Chair

Ad Hoc Committee on Election Procedures

Announcement:

The Donald E. Knuth Scholarship for 1992

The intent of the Donald E. Knuth Scholarship is to encourage the increase of knowledge about TEX and to sharpen the TEX skills of non-technical users.

One Knuth Scholarship will be awarded in 1992. The competition will be open to all 1992 TUG members holding support positions that are secretarial, clerical or editorial in nature, as determined by job title and duties, and not holding a degree with a major in a technical, scientific or mathematical subject area. The award will consist of an expenses-paid trip to the TUG annual meeting and to the Scholar’s choice from the short courses offered in conjunction with that meeting. A cap of $2,000 has been set for the award; however, registration fees for the meeting and short course will be waived, and not counted in the limit.

Schedule

March 23, 1992  Deadline for receipt of submissions
April 7–May 25, 1992 Judging period
June 1, 1992  Notification of winner

An announcement in TUGboat 12, no. 4, pp. 563–564, contains a description of the material that must be submitted by candidates, which includes a TEX project of the candidate’s choice, to be judged by the Scholarship Committee, and certain personal information.

The Scholarship Committee consists of Chris Rowley, Open University, U.K. (Chair); Nico Poppelier, Elsevier Science Publishers, Holland; David Salomon, California State University, Northridge; and Linda Williams, University of Tennessee Space Institute.

A brochure containing the TUGboat article and additional information is available from the TUG Office. To obtain a copy, or to request instructions on e-mail submission, write to the address below, or send a message by e-mail to

TUG@Math.AMS.com with the subject Knuth Scholarship request.

TEX Users Group

Attn: Knuth Scholarship Competition 658 North Main Street P. O Box 9506 Providence, RI 02940-9506 U.S.A.

What is the L^{A}TEX3 Project?

The accompanying description of the L^{A}TEX3 project from two of the active group members should help to explain what the project is about, and why there was an item on the membership renewal soliciting contributions. Our original intention was to provide background to the project in TEX and TUG NEWS late last year, but the decision to conserve TUG’s resources by postponing that edition led to the membership renewal being seen long before the explanation.

It is important to realize that TUG’s involvement is not that of directly providing financial resources, but rather of helping to attract support from a wide range of potential sources. Details of expenditure by the project team will be published in TEX and TUG NEWS. One of TUG’s major roles will be to make information about the project widely available. This will work at a number of different levels. For example, I have appointed Alan Hoenig as Board liaison with the project — he will be the TUG channel aware of the current status of the project, and able to summarize that information for us all, through the medium of TEX and TUG NEWS. We will also be publicizing and supporting the tutorials and workshops envisaged by the project, both to discuss the problems uncovered, and to allow the future users of L^{A}TEX3 the opportunity to see where current thinking lies. We may also expect to see more substantial information about aspects of the programme published in TUGboat; for example, Frank’s ‘Stockholm’ paper provides some helpful insights into the project and will appear sometime this year.

Much rides on the success of this project. It is the first time that TUG has contributed to a major project in this way — and L^{A}TEX3 is undoubtedly a major project. The widespread use of L^{A}TEX as the TEX tool of choice gives this particular project relevance to us all. The lessons learned will influence the way we participate in other areas which we believe will secure the future of (L^{A})TEX.

Among the current sponsors of the project, who have provided cash or some other support, are: Addison-Wesley, the American Mathematical Society, Arbor-Text, Aston University, Blue Sky Research, Digital Equipment Corporation, EDS Electronic Data Systems, Elsevier Science Publishers, Nordic TEX Users Group, the Open University, Personal TEX Inc, the Royal Institute of Technology (Stockholm), the Royal Military College of Science, Southampton University, the TEX Users Group, TEX88, TEXpert systems, the UK TEX Users’ Group, and Universitätsrechenzentrum Heidelberg. Individual members of TUG have contributed just under $1,000 (so far).

Malcolm Clark

President, TEX Users Group

The L^{A}TEX3 Project Fund

Some of you may be wondering why your TUG membership renewal form asked for contributions to L^{A}TEX3. The immediate reason is that the Board of Directors has agreed to set up and administer a L^{A}TEX3 Project Fund, to publicize this fund and to help in attracting wide-ranging support for it — from corporations, trusts and other external sources — and also from TUG’s institutional and individual members (that’s you!).

But before you reflect on how you can help, you would probably like to know a bit more about the project.

Brief history

The decision to produce an improved and expanded version of L^{A}TEX was taken during the 1989 TUG conference at Stanford. It was the result of two meetings between Leslie Lamport and groups of L^{A}TEX users which were followed by detailed discussions between Frank Mittelbach and Leslie Lamport.

Wider discussions have taken place about the various requirements for the new version, and a small group has produced prototype implementations of some parts of this system. The project’s technical director is Frank Mittelbach and its management is shared by him with Chris Rowley and Rainer Schöpf.

Aims

The L^{A}TEX3 project will produce a public-domain document preparation system which combines the power and excellence of TEX’s typesetting capabilities with ease of use for authors, editors and designers.

The most important principle, which is followed by the present L^{A}TEX and which will guide the new project, is the separation of generic mark-up from visual formatting.

A major new feature of L^{A}TEX3 will be a good interface through which designers can specify how classes of documents will be formatted and hence set up the style files which enable TEX to typeset documents according to these specifications. It will enable them to specify any font families and give access to a large range of typographic possibilities and to many detailed enhancements and extensions of the current facilities.

Thus L^{A}TEX3 will be able to process a far wider range of documents. It will also have a more robust and more flexible author interface (no fragile commands), with interactive error-handling linked to a help system. It will, of course, be multi-lingual.

The code will consist of modules with precisely specified interfaces, thus making future maintenance and extensions straightforward. Full documentation of all parts of the system will be produced as part of the project. Easy and rapid conversion to the new version will be ensured by wide dissemination of the information and utilities needed for up-grading.

The TUG community, especially providers of software (both public domain and commercial), will be kept fully informed about the progress of the project and its implications.

Schedule and finance

Work on the project started some time ago and it is scheduled to be completed by mid-1993. Since none of the volunteers is working full-time on this project, maintenance of this schedule is to some extent dependent on other demands on their time. It has also become clear that progress, which has been good so far, will depend on adequate financial support being available for the following purposes:

• enhancement of computer equipment and software for the core development team;

• purchase of books on typography and other related subjects;

• essential expenses (travel, accommodation, etc.) for occasional meetings of the project’s core development team; and also for meetings with others, for example: testers of early versions; publishers; typographic designers; suppliers of related software, etc.

So, now you know how your money will be used. Even if you have already returned your membership renewal, it is not too late to send your contribution to the TUG office, clearly marked L^{A}TEX3 — and many thanks, especially to all those who have already contributed.

And don’t forget that “every penny counts” — it does not matter how small your contribution is. To put this in perspective, if every member of TUG contributed just $10 then we should have an excellent basis from which to attract further funds from other sources.

Frank Mittelbach Chris Rowley

Scientific Publishing in Mexico

In this report, we want to give a picture of the TEX users’ community in Mexico, a bit of history and pride, and a brief description of the capabilities and needs of Mexican scientific printers. In view of the growing free trade atmosphere, this information may be useful for technical text publishing companies, TEX enterprises advertising in TUG publications — and freelance hackers who would like to live in a milder climate.

The transfer of knowledge

You can probably guess that the first printing shop in America was south of the border. Early colonial Mexico City got its first press in 1539 thanks to the efforts of the Friar and Bishop Juan de Zumárraga, who inspired the first Viceroy, Antonio de Mendoza, on the benefits of modern “high technology.” Giovanni Paoli, a master of repute, came to live in a milder climate, changing his name to Juan Pablos, First Printer of the Americas. His shop was at the Casa de las Campanas, somewhere downtown. An immediate benefit was a steady production of fine theology texts and prayer books; the longer range benefits included the growth and reproduction of Spanish culture and academic life, even during the stagnation of the seventeenth century. This and similar technical knowledge permitted the Bourbon kings to launch their economic reforms, and later channeled the intense literary life and political polemics of the last century.

Although Pablos’ shop preceded the one in Cambridge by a hundred years, newspaper printing technology had to be imported anew at the turn of the century. It quickly caught on, however, and was used with gusto when politics turned violent again, as were other gadgets such as railrods, telephones, and airplaines (Carranza was the first acting President to fly, in 1917). This pattern of technological spread has also applied to the particularly rapid computer revolution that started in Mexico in the early 60s: it spread via researchers, mostly from public universities, who belonged to formal or informal international scientific networks, and were thus informed and abreast of new developments. As in other scientific/technological areas, the pattern seems to be that these individuals can rather quickly reproduce prototypes in their home Mexican institutions. There sometimes the prototypes linger as a curiosity, a prized tool for academic research, but rather slowly and rarely diffused to industry.

Max Díaz brought the first TEX’80 version to the Instituto de Investigaciones en Matemáticas Aplicadas y en Sistemas (IIMAS), of the National University of Mexico (UNAM). It lay there as a curiosity in the Foonly mini-computer for a few months, while Max worked on View and a laser printer was found.

I had always had a typographic interest, so for the School and Workshop on Nonlinear Phenomena, which we had organized in Oaxtepec during December 1982, we had the possibility to test TEX on stationery and programs for the event, and the nerve to propose that the Proceedings volume should thus be nicely printed. (All the great gurus at the meeting were happy and willing to write hefty contributions.) Its 500-odd pages were completed in July 1983 with the part-time help of five science students who did the typing (Wolf, ed., 1983). We know that Donald Knuth’s TEXbook (1984) and Michael Spivak’s Joy of TeX (1986) were the first and second books in TEX and also that ours was the first with Springer. Unless readers correct me, third place does not seem too bad …

Good! So, why not write some more books!? At IIMAS we edited an engineering textbook on integral transforms (Márquez and de la Lanza, 1984), and one on quantum mechanics (de Llano, 1985): both turned out to be complete bureaucratic disasters (the floppies are still kept in a drawer). The National University of Mexico is the largest publishing house in Latin America, but back in 1985 it was impossible to run a profitable enterprise with exotic technology, let alone pay a fee per page. The editors of Springer Verlag were also discouraged. In retrospect, however, the effort was not all wasted (Wolf et al., 1986). Among the few people who have become TEX professionals in Mexico are some of those very students who typed those first texts. Also, as more researchers drooled with envy at the impeccably printed preprints (e.g., Wolf, Hlavatý, and Steinberg, 1982; Boyer, 1982), theses (Navarro Saad, 1985), and books (Sánchez-Mondragón and Wolf, eds., 1986; Wolf, ed., 1989) from IIMAS, the TEX virus continued to extend to other institutes and universities. Most important (and much more effective) was the establishment of a soft- and hardware enterprise by Armando Jinich and Max Díaz (both from IIMAS) that sold protected PCTEX and initially also provided maintenance and TEX training for personnel.

El Grupo de Usuarios de TEX

In December 1986 we convened La Primera Reunión del Grupo de Usuarios de TEX in Cuernavaca, a beautiful place 80 km south of Mexico City (Wolf, 1987). There were 57 participants representing a healthy variety of academic and governmental institutions, and publishing companies. The following year we counted 90 volumes produced in TEX, including the Enciclopedia de México.^{1} In successive years, users and interest have diversified: since 1988, the Presidential Address to the Nation read superbly; non-trivial ampersand tabulation macros have been written, although their commercialization is still hampered by antiquated software copyright laws; high-tech art and sui generis-formatting books on musical theory and critique (Estrada, 1990). Several physics, mathematics, and other academic journals are now printed in TEX: the Boletín de la Sociedad Mexicana de Física and Revista Mexicana de Física (published by the Mexican Physical Society); the Boletín de la Sociedad Matemática Mexicana and Anales del Instituto de Matemáticas; and institutional journals such as Avance y Perspectiva, Contactos, and Ciencia — Revista de la Academia de la Investigación Científica.

1[For more details on this and other projects in Mexico, see TUGboat 10 (1989), No. 4, pages 579–593 –Ed.]

Published since 1954, the Revista Mexicana de Física is the oldest and bestknown Latin American physics journal. Since 1987 (vol. 33), it has been produced with TEX with Fís. Arturo Sánchez y Gándara as TEXnical editor. Volumes 33 and 34 have contained an average of 15 articles and 174 pages per quarter (mostly in English, but Spanish and Portuguese are also accepted), or roughly 700 pages per volume (Sánchez y Gándara, 1990).

In January 1990, the editors of these scientific journals held a Round Table at the Fourth Grupo de Usuarios meeting. It was seen that desktop publishing units had been successfully established (with editors, typists, secretaries) at each of the institutions, but that each was devoted to one or two publications. Springer Verlag at Heidelberg, by contrast, manages 250 journals, handling everything beyond the “… we accept your paper … ” letter of the journal editor.

Clearly it was time for a publishing company to appear to handle these four or six scientific journals and bulletins, thus offering savings to the government subsidies that have nurtured them all. Easier said than done, but the 1991 Round Table at the Fifth MexTEX (entitled MERCADOTEXNIA) gathered some four TEX entrepreneurs who now carry on the production (although not yet the distribution) of one or two of these journals. At the same time, we felt the need for a software company to provide services and training for TEX users (Aurión Tecnología gradually bowed out of these activities during 1990). We now depend on a spidery web of altruistic pirates to provide us with new fonts (Cyrillic, Arabic, Hebrew), figure caretakers, global formats and general updating, an information clearinghouse and friendly distributors of hackers’ macros.

In 1989, the Centro Internacional de Física y Matemáticas Aplicadas (CIFMA AC, Cuernavaca) obtained a grant to further editorial competence^{2} by sending a couple of people to TUG annual meetings and to Springer Verlag in Heidelberg for training in journal management. Although Mexico’s human resources for scientific publications in TEX are good by Latin American standards (on a par with Brazil, and not too far behind Spain), we have had few links with international TEX users groups, commercial or academic. It was thus good to have with us at the 1991 meeting Ms. Betsy Dale of ArborText, to learn of the services provided by her and other American TEX companies. The Grupo de Usuarios de TEX is still an informal organization; its creation under the legal title of an Asociación Civil has been an idea raised at the yearly meetings and understood to be a hot potato with lots of work and little academic recognition.

2Consejo Nacional de Ciencia y Tecnología, Tipografía Científica Automatizada, P107 CCOT89/4517.

Some conclusions

It has also become clearer what kind of work TEX is good for, and what other choices are better for page design, legal bulletins, and pulp magazines. The consensus is that most ordinary text is easier and cheaper to produce with what’sits-name systems usually friendlier to secretaries and typo workers. Scientific text is TEX’s home territory, no doubt. Technical text edition, such as engineering journals and bulletins, reports and projects, is the gray area where TEX companies have to prove their cost effectiveness and accede to this considerable market. The Instituto de Investigaciones Eléctricas de Cuernavaca (largest engineering research institute in the country) has such an organization to handle its output. The Instituto Mexicano del Petróleo (still larger) is in the first stages of adopting the system for some of its publications. Probably the most compact expert group in the edition of scientific text in Mexico is at the Sociedad Mexicana de Física.^{3}

The prices per page of journal text quoted at this year’s meeting were comparable to those advertised in TUG correspondence and in TUGboat. Since salaries are considerably lower in Mexico, one may wonder why scientific printing can’t also be lower. It’s no great secret, not for general reasons but rather, for particular reasons. For example, at banks, one has to stand in line for an incredible length of time to deposit a cheque; appointments are not quite so binding as in the time-poor First World; one visit is seldom enough for most matters; federal mail takes at least a week in spite of ZIP codes; pulse telephone lines are still prevalent and they are noisy, so faxes easily fail; few personal computers are linked to networks and one may be left netless for days on end; as well, the local tax office increasingly over-administrates.

Things are bound to change, we are told, so text and printing services should become competitive in Mexico in the near future. Communications are being upgraded (and privatized), albeit slowly; technological parks are being set up near the border, in Querétaro. These parks provide company-incubating services to guide newcomers through the red tape of the establishment, as well as some ready communication lines. Regarding personnel, local expertise in handling scientific literature is ample but diffuse. It exists in students and graduates of the School of Science at UNAM and some five other universities in the country; there are many key-happy academics (who need, rather than sell, expert services — there is too much improvisation by otherwise over-qualified people), and some smart bilingual secretaries. Probably some 200 individuals in all. Lastly, the professional profile of a scientific/technical editor is only beginning to be appreciated.

There is a long tradition of book production in Mexican art and literature publishing houses. The major Spanish, European and American book companies are represented and routinely translate and print technical books. For all these pluses, scientific literature is not yet handled in a fully professional and efficient way; there is more than one niche open for enterprising people.

3Sociedad Mexicana de Física, AC, Apartado Postal 70-348, 04510 México D.F., contact Arturo Sánchez y Gándara and José Luis Olivares.

References

Boyer, C.P. The Geometry of Self-Dual Einstein Spaces, IIMAX-CT Notas, No.

23. México: UNAM, 1982. Estrada, J. Los Sonidos en Juan Rulfo. México: UNAM, 1990. de Llano, M. Mecánica Cuántica. México: Facultad de Ciencias, UNAM, 1985

(unpublished). Márquez, P. García and C. de la Lanza. Ecuaciones Diferenciales y en Diferencias.

México: Facultad de Ingeniería, UNAM, 1984 (unpublished). Navarro Saad, M. “Métodos de teoría de grupos en el estudio de aberraciones de

tercer orden.” B.Sc. thesis. México: Facultad de Ciencias, UNAM, 1985. Sánchez y Gándara, A. “Aplicaciones de la tipografía automatizada al manejo de

información científica.” B.Sc. thesis. México: Facultad de Ciencias, UNAM, 1990.

Sánchez-Mondragón, J., and K.B. Wolf, eds. Lie Methods in Optics, Lecture

Notes in Physics, Volume 250. Heidelberg: Springer Verlag, 1986. Wolf, K.B. “News of TEX Users in Mexico.” TUGboat 8 (1987), no. 1, page 7. —— , ed. Nonlinear Phenomena, Lecture Notes in Physics, Volume 189. Heidel-

berg: Springer Verlag, 1983. —— , ed. Lie Methods in Optics, vol. II, Lecture Notes in Physics, Volume 352.

Heidelberg: Springer Verlag, 1989. —— et al. Manual de Lenguaje y Tipografía Científica en Castellano. México:

Editorial Trillas, 1986. Wolf, K.B., L. Hlavatý and S. Steinberg. Nonlinear Differential Equations as

Invariants under Group Action on Coset Bundles. I. Burgers and Korteweg-de Vries Equation Families. Comunicaciones Técnicas IIMAS, No. 219. México: UNAM, 1982.

Kurt Bernardo Wolf

Instituto de Investigaciones en Matemáticas Aplicadas y en Sistemas, Universidad Nacional Autónoma de México

Cuernavaca

Typographer’s Inn

In the desktop publishing (dtp) world at large there seems to be an upsurge in interest in typography, now that the first flush of enthusiasm among wysiwyg devotees is beginning to settle.

Ever anxious to be the first kid on the block, ttn is proud to introduce this column as a forum for discussing matters typographic. Here you can ask for help, grouse about the inadequacies of typefaces, laugh over some typographic gaffe, or shout about the new things you’ve done or seen done.

As a new column, we have no contributions yet, so I’m going to stand on my soapbox and shoot my mouth off, in the hope that it will stimulate you to write and fuel the debate.

First off is the question of training: how do we persuade new users (principally of tex, but actually of almost any word processing or dtp system) that spreading fonts across the page like peanut butter across hot toast is not necessarily the route to typographic excellence? I’m biased—I run typography training courses (on demand, for cash), so I try to explain that communication of ideas is the principal task of the typographic designer, and I try to equip new users with some basic skills in layout and font selection to help them do this. But I’m sure there are other ways, and I would like to hear about them.

That raises the second point: the Computer Modern fonts. They are ‘standard’, that is, they come shipped with almost every version of tex, so that a document using this known set of font files is portable worldwide. This is a huge advantage on the one hand, and a huge drawback on the other, because so many reviewers/commenters on tex gripe that “it only has one typeface”. Although this is rubbish (tex actually works with almost any typeface from any source, given the right conversion tools) it is not obvious to the beginner, or even the seasoned user, just how to go about getting and implementing new fonts. The revised font selection scheme for l^{A}tex will make life easier, but I for one haven’t had time even to look at it yet. Can someone perhaps summarize the steps for a new user of l^{A}tex?

Moving right along, there is the related question of layout. In l^{A}tex, the page layout is pre-defined, and it takes a fairly sophisticated hacker to change the design. In Plain tex, you have to do everything yourself from scratch. Much play is made in the documentation surrounding texware of the virtues of fully marked text, and its inherent portability, but we still need the middle course between Plain tex and l^{A}tex, which would provide the management macros such as sectioning, indexing, table-of-contents, cross-referencing and the handling of illustrations, without imposing a visual style which is hard to change. Anyone volunteer to write this? Or to tell me why it’s a silly idea?

What about a bibliography? I have references to (mainly) European books on typography, but we need a more worldwide spread for it to be useful. This could be stored on a listserv and on an ftp–accessible machine for people to retrieve. And what about starting up a TYPO-L list?

One of the reasons I use tex is because of the control it gives me. In most visual dtp systems, typographic quality takes a back seat to claimed ease of use. Just pick up almost any piece of dtp’d rubbish that pops into your mailbox and study it hard and critically: are the baselines straight? Is the interletter spacing correct? Are the margins evenly aligned, especially in itemised lists? Are the lines evenly spaced? All these things are taken for granted in tex, because it does them so well, but there are people out there unwittingly producing poor-quality work through no fault of their own except lack of knowledge.

While we’re on the subject, if tex is so hot at these things, why isn’t everyone using it? Some while back I was talking to the graphic designer at a nearby university where they use Ventura. Under his tutelage, they produce excellent work, but not all institutions are design-aware enough to employ someone as skilled as him to oversee the quality of publicly-consumed dtp. One thing he did mention, though, was having worked long and hard on a thesis which contained a lot of math. Yep, you guessed. They had to position every symbol in every equation by hand, with a mouse, because they were simply unaware that tex existed and could do the job automatically in a small fraction of the time.

Finally (for now), what about literacy? It was fashionable among ‘progressive’ educationalists and educators some while back to claim that a knowledge of punctuation, spelling and grammar was almost unnecessary these days. Funny how they’ve all gone quiet now that the labor market has tightened up and only the better-qualified applicants get jobs. But we still see ‘Street’ shortened to ‘St.’ when there is plenty of space for the whole word; indecision about punctuation after ‘Ms’, ‘Mr’ and the like; misplaced apostrophe’s; unnecessary hyphens, and a whole pile of little itty bitty things which may be unimportant in themselves, but contribute to the overall impression of sloppiness. Any compositor of the hot-metal generation would blush if accused of one of these slips, but now that anyone can do their own typesetting, how do we prevent the slide into the garbage can? Do we even need to worry about it?

   So tell me. Mail me (cbts8001@iruccvax.ucc.ie) with a Subject: line of

TTN Typography. Or mail the editor. And let’s make this column the place to discuss typography (it could give the phrase “letters to the editor” a whole new meaning).

Oh, yeah. A drink on me at the tug meeting in Oregon (1992) for the first mail saying what the significant typographic change in this article is.

Peter Flynn

University College Cork

New Publications

Claudio Beccari. L^{A}TEX — Guida a un sistema di editoria elettronica. Milano: Editore Ulrico Hoepli Milano, 1991. ix + 399 pp. ISBN 88-203-1931-4.

• The book is the first and only one in Italian at this moment. Although it contains an introductory chapter for the beginner, its main purpose is to give the reader advanced information on L^{A}TEX (and partially on TEX) in order to prepare sophisticated macros, style files, and to make clever use of fonts.

Arvind Borde. TEX by Example. New York: Academic Press, 1991. 169 pp. (softcover), US$19.95. $18 for TUG members. ISBN 0-12-117650-9.

• This book teaches entirely by example. Input and output are shown sideby-side, allowing users to spot quickly how to get symbols, how to obtain italics, how to use bold type, how to start new paragraphs, and so on. The alphabetic appendix is both an index and a glossary, which systematically lists all commonly used features of TEX.

Victor Eijkhout. TEX by Topic: A TEXnician’s Reference. Wokingham, UK: Addison-Wesley, 1992. viii + 307 pp. (softcover), US$30.00. $27 for TUG members. ISBN 0-201-56882-9.

• TEX by Topic is a systematic reference manual designed as a companion to tutorial guides about TEX. Not for the novice, it is meant for users with a basic understanding of TEX who have specific questions, or who are keen to explore the full potential of this unique computer typesetting system.

Jane Hahn. L^{A}TEX for Everyone. Mill Valley: Personal TEX Inc., 1991. xi + 346 pp. (softcover), US$19.95. $18 for TUG members. ISBN 0-9631044-0-3.

• L^{A}TEX for Everyone is a tutorial as well as a reference guide. Written with the beginning user in mind, it is designed to be easy to read and understand, and contains examples of all the basic commands. The experienced L^{A}TEX user, however, will also find this book to be a valuable reference tool.

Raymond Seroul and Silvio Levy. A Beginner’s Book of TEX. New York: Springer Verlag, 1991. 283 pp. (softcover), US$29.95, DM58,– . $27 for TUG members. ISBN 0-387-97562-4.

• Seroul and Levy’s book is an excellent beginner’s text. It is meant for careful reading, in which case it gives a surprising amount of detail. The book also has a reference section. This is an expanded translation of Seroul’s book, Le petit Livre de TEX (Paris: InterEditions, 1989).

“Hey — it works!”

The Cauchy Integral

When the applied sciences use complex numbers they generally need a symbol for the Cauchy integral, which is a regular integral sign with a bar across. The following macro will do:

   \def\intcauchy{\mskip3mu-\mskip-18mu\int}

and works pretty well with most L^{A}TEX size options, although the best is as always the 10pt default size. If you do not like \def you can use the standard L^{A}TEX commands:

   \newcommand{\intcauchy}{\mskip3mu-\mskip-18mu\int}

Now you can write in display math:

− ∫_{∞}^{∞} f(z) dz = lim_{R→∞} ∫_{R}^{R} f(z) dz

by using the new command \intcauchy as you do with regular integral signs.

By the way, ISO standards require the differential sign to be typeset with Roman type, a rule that is very seldom observed even in the most qualified mathematical papers.

Claudio Beccari

Politecnico di Torino, Italy

Dealing with Very Long Paragraphs

The following problem would never have occurred to me — I am a mathematician, and I do not write really long paragraphs. But habits are different, and so it happened that the other day a mathematics historian I know mentioned the following problem.

He complained that, once in a while, TEX reported insufficient memory when compiling some unusually long paragraph. Of course, he admitted that it would be possible to give in and introduce an additional paragraph. But certainly that would not be completely satisfactory. Another way out would be to use some bigTEX-Version — neither elegant, nor convenient.

Fortunately, there is yet another solution.

If a paragraph is too long for TEX’s memory, then certainly it should also be long enough to prescribe the \parfillskip! Therefore the following definition introduces an “invisible paragraph” \nopar. It can be inserted into a long paragraph — for instance, just before the point where TEX’s memory was exhausted — or even just “for safety” when a paragraph seems to be too long. But beware: when inserted too early \nopar may produce a nasty underfull \hbox!

   \def\nopar{{\parfillskip0pt\parskip0pt\par\noindent}}

Note the double pair of brackets — they serve to restore the previous settings of \parskip and \parfillskip after the invisible \par. See also The TEXbook, p. 100 for a description and dangerous bend problem 14.15, and p. 315 for the solution.

Peter Schmitt

Universität Wien, Austria

                                          a8131dal@awiuni11.bitnet

(L^{A})TEX News

IL^{A}TEX

As of December 18, 1991, International L^{A}TEX (or IL^{A}TEX for short) does not exist any more. If it is found on any TEX servers, it is outdated and should be deleted. The latest version of L^{A}TEX, dated December 12, 1991, has made this possible, since it incorporates the functionality of IL^{A}TEX. Please use this (or a newer) version.

The following parts were not incorporated into L^{A}TEX:

The Generic Local Guide. It will be released as a separate distribution at a later time. I have a lot of responses to check first, though.

The support for multi-lingual documents. The Babel system by Johannes Braahms provide an alternative possibility. If you prefer my system, it can still be fetched via anonymous ftp from ftp.th-darmstadt.de, directory pub/tex/initex, files mlinitex*. I think that parts of this system might influence the Babel development in the future.

I hope that I have served the TEX community with the IL^{A}TEX support and look forward doing new, other, work in the TEX arena.

Joachim Schrod Darmstadt

                             schrod@iti.informatik.th-darmstadt.de

News from Around …

CyrTUG Irina Makhovaya

Mir Publishers, Moscow

I’d like to send you CyrTUG’s warmest regards. CyrTUG is a new group, uniting TEX users in Russia. Our objectives are to further development of TEXnology in Russia, coordinate developments and standardization, create a Cyrillic TEX, and support contacts with TEX users from the USA, England, France, Germany, Holland and other countries. We understand that there is serious interest in a Cyrillic TEX in the United States and other countries. We have also heard about a Slovenian Users Group.

Our constituent meeting took place at Mir Publishers, in Moscow, May 23– 24 of this year; coincidentally, this was also the Cyrill and Mefodiy days — a celebration of the Slovenian languages. In all, twenty-three representatives from twelve establishments in four Russian cities attended the meeting. The number of TEX users at each of these institutions ranges from one to 100 people, so that effectively some 250 users from across Russia were represented at the meeting.

The following individuals were elected to the Board of Directors:

Joseph Romanovsky, President, LSU, Leningrad

Irina Makhovaya, Executive Director, Mir, Moscow

Vladislav Andrushchenko, IRL, Moscow

Mark Doubson, CEMI, Moscow

Eugene Pankratiev, MSU, Moscow

Alexander Samarin, IHEP, Protvino

Anatoly Urvantsev, NSU, Novosibirsk

A second meeting took place at the Institute of High Energy Physics (IHEP), in Protvino, a suburb of Moscow, August 22–23. In September, some reports on our activities were presented at the EuroTEX conference in France, and some will appear in TUGboat. We appreciate very much the invitations from EuroTEX to our colleagues.

We plan to hold our annual meetings in May. The first Annual Meeting of CyrTUG will take place at IHEP, Protvino, in 1992. We would be very glad to see TEX gurus, TEXperts, and anybody from TUG, the UKTEXUG, DANTE, GUTenberg, NTG, and other national groups. We would like to invite everyone who is interested in TEX, Russia and international cooperation.

SibTUG Dr. Anatoly L. Urvantsev

Computing Center, Novosibirsk University

This is official information about the founding of a TEX Users Group in Siberia. The inaugural meeting took place 6 Dec. 1991 at Novosibirsk State University. Fifteen people, representing ten organizations from the Siberian cities of Novosibirsk, Krasnojarsk and Barnaul, took part.

Members of the board of SibTUG include:

Dr. Anatoly L. Urvantsev (Computing Center, Novosibirsk) – Chairman Dr. Vladimir N. Djatlov (Novosibirsk University) – Executive Director Il’a R. Roussakov (NAUKA Publishing House, Novosibirsk) Prof. Anatoly M. Fedotov (Computing Center, Krasnojarsk) Dr. Gennady I. Algazin (Altaj University, Barnaul) Prof. Vladimir A. Vasilenko (Dialogue-Siberia) – Financial Inspector

There was strong interest from scientists and other specialists from other cities, including Omsk, Tomsk, Tjumen, Kemerovo, and Irkutsk to form local branches of SibTUG. The Siberian division of the Academy of Sciences comprises some 80 academic institutes of various scientific areas of interest (about 40 are in Novosibirsk). Many other organizations, such as publishing and printing houses, software and hardware firms, and so on from this region are interested in highquality publication of scientific, educational and engineering literature.

The Siberian Science and Education system has good connections with colleagues and centers in different countries. We also enjoy good cooperation with scientists from other republics of the former USSR, especially the Asian republics. As a result, we are not only focussing on Cyrillic versions of TEX, but also on a whole rainbow of TEX possibilities.

Because SibTUG is, naturally enough, the Siberian part of CyrTUG, we shall coordinate our efforts. Nevertheless, given our large region and poor connections (transport, mail, e-mail, etc.) with Moscow, we ask the TEX world to contact us directly.

From our side we shall support the TEX community in your efforts to extend the use of TEX and to include high-level software specialists in the process of producing new TEX possibilities.

The main problems for SibTUG and TEX users in Siberia are:

– the absence (in practice) of books on TEX

– the absence of good versions of TEX

– the absence of TEX materials

– poor communication between users

– limited knowledge about happenings in the larger TEX community

We shall concentrate our efforts on solving these problems. We also want to give our users the opportunity to use Russian language TEX materials as well as materials on other languages of Siberian countries. So, our first steps will be to organize:

– a library of TEX materials (books, proceedings, magazines, etc.)

– a collection of TEX software products and demonstrate their capabilities – courses for different levels of users in Novosibirsk and other Siberian cities – better communications between TEX users inside our region and beyond

However, problems such as exchange of the rouble and other national concerns are currently hampering our efforts to acquire these materials in legal ways. We greatly appreciate any support and help from all sides as we address our primary needs, especially during our start-up period. In spite of these difficulties, we are determined to solve these problems, and offer our colleagues in other TEX users groups to cooperate with us. For our part, we shall be very glad to return that help to the TEX community also.

On behalf of SibTUG,

Dr. Anatoly L. Urvantsev Chairman, SibTUG

Computing Center  Tel (3832) 350-454
prosp. Lavrent’eva 6  FAX (3832) 324-259
Russia, 630090, Novosibirsk E-mail uran@isi.itfs.nsk.su

Reports on Meetings

TUG ’91: 15–18 July 1991 Dedham, Mass.

The 1991 Annual Meeting of the TEX Users Group took place at the Dedham Hilton, just outside of Boston, Massachusetts; the theme was “TEX Inroads into Publishing: Document Design, Production and Interchange.” Judging by the number and diversity of presentations and attendees, this was a truly successful meeting. And as with all TUG meetings, there was a lot more than just publishing being discussed: new font and graphics packages and applications, and future directions of TEX were also very much on people’s minds.

Each day was dedicated to a specific sub-theme, which made it easier for attendees to focus on their preferred areas of specialisation. In addition to some 193 people who registered for all four days, there were 50 registrants who came specifically for the Monday program, 11 who came on Tuesday, and 6 on Wednesday.

The meeting was launched with a provocative keynote address by Nico Poppelier of Elsevier Science Publishers (Holland). In all, there were 28 presentations, 4 panel discussions, and 7 workshops, with the TUG Business Meeting, the Questions and Answers session, and an Historical Round Table also vying for everyone’s attention.

In 1992, the TUG Annual Meeting will be held in Portland, Oregon; Program Coordinator will be Mimi Lafrenz, with Doug Henderson, Helen Gibson and Ron Whitney on the Program Committee. After four years of working on TUG’s Annual Meetings, I must confess to feeling a great sense of relief at handing over the task to a new Coordinator — and I look forward to attending this year’s meeting as a regular TUG member. See you all there!

Christina Thiele

TUG’91 Program Coordinator

CsTUG: 18–20 Sept. 1991 Skalský Dvůr, Czech.

The 1991 meeting of the Czechoslovakian TEX users took place in Skalský Dvůr, located to the southwest of Prague, surrounded by hills and forests, far away from the bustle of the big cities. It’s just the right place to concentrate either on a typesetting system or on the outstanding Czechoslovakian beer.

The TEX-related portion (from September 18–20) of the Ninth Summer School on Computing Techniques in Physics, in which the meeting traditionally is embedded, presented papers on: new features of TEX v. 3, SGML in the publishing world and other SGML-related topics, L^{A}TEX and related software, TEX and graphics, exchange of knowledge, experience, software within the TEX environment, and working with .

There were about 120 attendees, from Bulgaria, Finland, Germany, Hungary, Libya, Poland, Romania, Switzerland, USA, USSR (at that time)^{1} and the UK. Because of the international audience the closing discussion about the needs in different Eastern countries was a very important topic. The people who had the problems and those who had the solutions were in the same room at the same time and were talking together. So the discussion was a great success.

But one example of the importance of information in the Eastern European countries should be mentioned: the universities in Romania were forced to buy a commercial DTP system because TEX was unknown there! And TEX was what they were looking for.

Friedhelm Sowa DANTE Treasurer

EuroTEX91: 23–25 Sept. 1991 Paris

The 6th European TEX Conference was held in Paris’ Latin Quarter last fall (Sept. 23–25) at the Maison de la Mutualité, followed by a one-day GUTenberg Conference.^{1} Highlights from the two conferences included four 4-day courses before the week of the meeting (a special mention should go to Philip Taylor and his course on advanced TEX and TEX macros), some two dozen presentations with simultaneous English-French translation, two panels, commercial sessions, a question period and parallel Birds-of-a-Feather sessions. There was a large exhibit area, and numerous terminals or stations for attendees to access the Internet.

1Tempus fugit.

The social activities included a film shown on the Géode’s 1,000 m^{2} hemispheric screen, dinner inside the Cité des Sciences and a bus tour of Paris by night.

And yet … these attractions weren’t enough to convince people to come; perhaps it wasn’t a good week, perhaps the fees were too high, perhaps there are now too many TEX meetings around the world … There were a little over one hundred attendees (coming from 21 nations), although we had planned for twice as many (based on attendance figures from previous GUTenberg annual meetings). In all, there was too much effort for too little return, and the GUTenberg Organizing and Program Committees are a little disappointed. All the ingredients for a successful meeting were present but …

The presentations have been published in a double-issue (nos. 10–11) of the Cahiers GUTenberg; however, the two panels (TEX in Europe, and L^{A}TEX V3) and the BoFs were not included.

In conclusion, as a (L^{A})TEX user, I can say that I was satisfied; however, as manager of the local users group, I was also quite disappointed.

Bernard Gaulle GUTenberg President

TUG VP for French-Speaking Users

TECH’92: 16–18 June 1992 Les Diablerets, Switz.

Call for Papers — Deadline Extended!

The 1992 GUTenberg meeting will be held in Les Diablerets, Switzerland and will be organized by the Observatory of Geneva.

Deadline for the Call for Papers has been extended to 15 April 1992. The conference will address the problems of TEX’s social environment at the support and training levels, and the use of TEX in non-scientific applications. Of course, more general TEX and DTP-related subjects are also welcome. Send submissions

and inquiries to megevand@scsun.unige.ch

Denis Mégevand

Président du comité d’orgaanisation

1Thanks to TUG, UKTEXUG and GUTenberg for their financial help in sponsoring Eastern representatives to attend.

TUG’92: 26–30 July 1992 Portland, Oregon

TEX in Context: Resources, Support Tools, and Comparative Studies

The thirteenth annual conference of the TEX Users Group will be held in Portland, Oregon, USA. This is going to be an exciting week of lively discussion and informative presentations covering a broad range of subjects relating to the various uses of TEX throughout the world.

The developing technologies that supplement the power of TEX — fonts, graphics, format macros, and archives — will be explored in presentations followed by question and answer sessions. Ever-popular workshops will be run several times each, to ensure the possibility of participation.

Panel discussions following themes of the presentations will address fonts, graphics, and include a special discussion on archives and the tremendous information available. The current 1991 Resource Directory will be just one of the many tools explored and explained.

The recently restored Benson Hotel in downtown Portland has agreed to host the conference and meetings at a rate close to 50% of their regular price. The Benson was the only hotel willing to allow more than one person in each room at the same base price. So don’t be overcome by the beauty of this old charmer; it really was the best deal we could muster. If you want accommodations for less than the $89 standard rate, they will be available, in limited supply, at a nearby inn. The Program Committee is looking into local universities for possible housing.

A new feature this year is the all-volunteer squad of assistants ready to help with information on anything about the program. If you would like to help with this year’s program, please indicate your ideas on the conference registration form.

For families and friends, we will have some organized adventures running concurrently with the meetings. Doug Henderson is organizing a fund-raising bowling tournament. There will be many things to see and do in Portland. See the conference brochure for more details, due in your mailbox in February.

One last event that bears mentioning is the Mt. Hood Festival of Jazz, beginning Friday, July 31st and continuing through Sunday, August 2nd. Tickets will be made available through TUG. Please mark your registration form if you are interested.

Members of the Program Committee for TUG’92 include: Helen Gibson, Doug Henderson, and Ron Whitney.

Mimi Jett

TUG’92 Program Coordinator

TEX and TUG NEWS Vol. 1, No. 1, 1992 23

TUG Board Activities

Summary of Board Meeting^{1}

The TUG Board of Directors met for 2 days before the 1991 Annual Meeting, 13–14 July 1991, at Dedham, Massachusetts. Board members who were present all or part of the two days included Nelson Beebe, Barbara Beeton, Lance Carnes, Malcolm Clark, Allan Dyer, Bernard Gaulle, Regina Girouard, Roswitha Graham, Alan Hoenig, Hope Hamilton, Doug Henderson, Patrick Ion, David Kellerman, David Kratzer, Joachim Lammarsch, Kees van der Laan, Pierre MacKay, Craig Platt, and Christina Thiele. Also present was Ron Whitney, TUG’s Technical Director/Business Manager.

Board members not able to attend were: Bart Childs, John Crawford, David Fuchs, and Dean Guenther.

During this meeting, some 43 motions were discussed and voted upon; the main issues are presented here. In keeping with a decision made at the March board meeting [summary in TTN 0,0], we are publishing this summary of the meeting minutes for the membership of TUG so that you may be aware of the work being done by TUG.

Malcolm Clark was elected Interim President, to serve from January 1, 1992 through the end of 1992. The Board elected the following from its members to serve as interim officers: Christina Thiele as Interim Vice-President and Secretary, and Allan Dyer as Interim Treasurer. The latter two officers will serve until replacement appointments are made by the newly elected board.

The President, Nelson Beebe, was directed to make an offer of employment to Ron Whitney, to provide for a formal agreement between Ron and TUG.

The need for a Technical Council has been apparent for a few years now and the Long Range Planning Committee (David Kellerman, Chair) was instructed to formulate the initial charter for the Technical Council.

Malcolm Clark believed the need for a specific European Coordinator is no longer there, so the Board voted to abolish the position that Malcolm has held, and added a note of appreciation for the work that he has done.

A budget for 1992 had not be developed so the Executive Committee was charged with coming up with one in the months following the July board meeting. A number of budget-related items were discussed.

• In an effort to better distribute funds, the TUG Office will modify their procedures to move toward fund accounting procedures.

• It was agreed that a maximum of $40.00 should be allocated to each board member without e-mail access to gain that access, thereby allowing them to participate in board-related business.

• An adverse reserve fund (initial amount of $50,000) will be established for the specific purpose of funding TUG’s liabilities in the event of an unforeseen and major catastrophic financial event.

1[These reports were prepared with the assistance of David Kratzer and Regina Girouard –Ed.]

• A motion to cap travelling scholarships and contributions to other societies at $4,000 failed; a subsequent motion to include this amount in the 1992 budget was passed.

• Committees requiring funding must now submit a budget to the TUG Office, for approval by the board.

• An effort will be made to receive additional funds for certain projects, such as TEXhax, the L^{A}TEX3 project, and others. The TUG Office was directed to establish restricted funds for such projects.

The board adopted a policy stating that it was in favour of reciprocal membership arrangements between TUG and other user groups; interested groups should submit proposals to the TUG Office.

The board asserted its intentions to continue using terms associated with TEX and , as has been done for at least 14 years (e.g., dvi, tfm, TEX, etc.). The issue arose out of examples of the use by other groups of the letters “dvi” for either their group’s names or products. It was deemed necessary for TUG to declare the long-standing and specific use of “dvi” for “Device Independent”, and its integral use within the TEX-using community. On the related matter of copyright, it was decided that TUG would adopt an appropriate form of the new general public license as the copyright license for TUG documents.

A Site Selection Committee was named, to select the 1992 site and to draft guidelines for the selection of future sites. Members of the committee are: Bernard Gaulle, Mimi Lafrenz, Christina Thiele, and Ron Whitney.

Draft guidelines on the Knuth Scholarship were accepted; Barbara Beeton will serve as board liaison between the Knuth Scholarship Committee and the board. Chris Rowley, David Salomon, and Nico Poppelier were appointed as the core of the Knuth Scholarship Committee (Chris Rowley, Chair) for the period 1991–1992; Linda Williams, the most recent previous Scholar, was appointed for a term of one year.

To defend TEX against misrepresentation, the Board will offer a $16.00 reward for the first person pointing out a published possible misrepresentation of TEX and reporting it to the TUG Office.

The board approved the following policies on attendance at official meetings: all TUG members may attend board meetings; all board members may attend Executive Committee meetings.

The TUG Office has been authorized to maintain a file of certified statements containing a minimum set of information on various TEX implementations so that purchasers may verify that a particular TEX system has measured up to approved standards. Vendors will be encouraged to state in their ads that the certified statements are on record in the TUG office.

There was much discussion about the new election procedures which will cause all members of the Board of Directors to be elected by the TUG membership. The procedures outlined in TTN 0,0 were adopted; by the time you read this, the results will be known. [See Election Results elsewhere in this issue – Ed.]

The Board adopted the following Mission Statement for TUG, following lengthy discussions of a draft proposal from the Mission Statement Committee (Christina Thiele, Chair):

The TEX Users Group (TUG) provides leadership

1. to encourage and expand the use of TEX, , and related systems

2. to ensure the integrity and portability of TEX, , and related systems

3. to foster innovation in high-quality electronic document preparation

Since there is now an organized users group in Japan, the head of this new users group will be invited to join the Board as a special vice-president.

The positive response that the newsletter TEX and TUG NEWS has received prompted the Board to authorize $5,000 for another issue in 1991. [Due to TUG’s financial siuation it was later decided not to publish another issue in 1991 –Ed.]

Other issues discussed included: the role of board liaison, special vice-presidents, and “porte-parole” (spokesman for TUG), TUG’s non-profit tax status, TUG’s legal counsel, annual meetings outside of North America, responses to queries regarding TEX’s future, reports from various committees, reports from various user groups, operating procedures manual(s).

Summary of Executive Committee Meeting

The Executive Committee held an evening meeting on 17 July 1991, during the 1991 Annual Meeting, at Dedham, Mass. The Executive Committee is made up of: Nelson Beebe (President), Malcolm Clark (1992 President), Christina Thiele (Interim Vice-President and Secretary), Allan Dyer (Interim Treasurer), Alan Hoenig (invited by Nelson Beebe), and Ron Whitney (Technical Director/Business Manager), a non-voting member.

Others present included Doug Henderson (recording secretary), and several board members.

During this approximately 4-hour meeting, the following major agenda areas were discussed: short-term requirements and issues raised during the general meeting, reports from various committees, and the TUG budget.

Charlotte Laurendeau will be leaving the TUG office after several years of service to TUG; the Executive Committee approved the recognition of her service by giving her a gift during the closing session of the meeting.

Ron Whitney, who is TUG’s Technical Director/Business Manager, has been serving TUG since the departure of Ray Goucher but his role has not been defined clearly. Several issues were discussed and the Executive Committee gave Ron authority to deal with many items which came up during the Board meeting, the general meeting, and the Executive Committee meeting.

The Executive Committee will meet in the fall to finalize the budget and handle some of the items that were put on hold at this meeting.

In the future, when TUGboat is ready for distribution, the TUG Office will make shipments to trans-Atlantic destinations before domestic mailing.

Mimi Lafrenz was named the 1992 Program Coordinator. Mimi has also volunteered to devote a significant amount of work on the L^{A}TEX3 project.

The Publications Committee (Kees v.d. Laan, Chair) report included the following points: news items will be shifted from TUGboat to TEX and TUG NEWS, which will be published quarterly at 28 pages per issue (7 × 10in); TUGboat will be limited to 800 pages per year; the resource directory, which all TUG members received, will also become an item for individual purchase; a catalogue of TUG items for sale should be produced; a TUGboat index should be prepared (based on author and keyword).

The Membership Committee is considering a two-tier system for all new items where members would get better prices than non-members. Lacking a consensus, the Committee made no recommendation at this time. More effort will be put into retaining current and attracting new TUG members.

The Local User Group Committee (Malcolm Clark, Chair) will be presenting its final report after the current meeting. Two preliminary reports had already been submitted to the board.

The Site Selection Committee (Christina Thiele, Chair) selected Portland, Oregon as the site for the 1992 Annual Meeting, considering in part the following factors: proximity of a major airport, range of hotel choices, Univ. of Oregon computer labs close by, a known concentration of TEX users and companies. Another factor in the Committee’s selection was the desire to move the meeting to a different geographic region.

Marketing of TUG and its products was discussed. It was decided that a tentative budget allocation of $20,000 be included in the 1992 budget.

The Executive Committee directed Ron to look into reducing the prices for back issues of TUGboat. Currently, old issues of TUGboat are still available for $17.00 each. When new user groups start up, TUG materials will be made available at nominal costs.

The startup of a restricted fund for TEXhax was approved. This would see monies allocated directly towards TEXhax, both as a budget item and as a focus for contributions. The money so allocated may not be used for other purposes.

There was discussion of starting up a similar restricted fund for research into TEX extensions; the Executive Committee voted to postpone this move for future board discussion.

The Executive Committee discussed the need for the annual meeting to break even rather than lose money. There will be a concerted effort to achieve this goal with the 1992 meeting.

The newsletter, TEX and TUG NEWS, should be budgeted for 4 issues at $5,000 each in 1992. Sponsorship was also discussed, and there was a consensus that $500 be spent for design of the newsletter.

By consensus, it was decided that a total of $18,000 should be included in the 1992 budget for production of six issues of TEXniques.

Other issues discussed included two-tier membership, a subscription category for TUGboat, donations to TUG, producing an annual report, more input on the mission statement, having an open-ended business meeting, having the board reconvene after the business meeting, initiating a “reach-out” component for new attendees at TUG meetings.

Next Board Meetings

15–16 February 1992, Aston University, UK

25–26 July, at TUG’92, Portland, Oregon, USA

Upcoming Events

23 March TUG Knuth Scholarship. Deadline for receipt of submissions.

TUG Office

                           tug@math.ams.com

24–27 March dante conference in Hamburg.

Reinhard Zierke or Gerhard Friesland-Köpke

                           dante92@informatik.uni-
                           hamburg.de

7–10 April EP’92: “Electronic Publishing, Document Manipulation, and Typography.” Swiss Federal Institute of Technology, Lausanne.

EP’92 — Dép. d’Informatique, IN Ecublens, CH–1015 Lausanne, Switzerland

                           ep92@eldi.epfl.ch

Spring UK TEX Users’ Group: Tentative meeting scheduled Peter Abbott

                           p.abbott@aston.ac.uk

May CyrTUG: 1st Annual Meeting. Institute of High Energy Physics, Protvino (near Moscow).

Irina Makhovaya

                           irina@mir.msk.su

4 June NTG meeting: “Scientific Publishing and TEX.” CWI, Kruislaan 413, Postbus 4079, 1009 AB Amsterdam.

Gerard van Nes

                           vannes@ECN.NL

16–18 June TECH’92, a GUTenberg conference: “The Dark Side of TEX”/ “La face cachée de TEX.” Les Diablerets, Switzerland.

Denis Mégevand

                           megevand@scsun.unige.ch

June UK TEX Users’ Group: “Design Issues”: A visit to the Dept. of Typography, University of Reading.

Peter Abbott

                           p.abbott@aston.ac.uk

27–30 July TUG’92: “TEX in Context.” Portland, Oregon.

TUG Office

                           tug@math.ams.com

In Memoriam

Cathy Booth

Cathy Booth died on July 27th, 1991, of cancer of the kidneys. For the last year she fought the disease consistently, cheerfully and resolutely.

Cathy was born on June 19th, 1952. She attended Exeter University, where she completed a BSc in Geography, and in 1984, she began work there as Computing Development Officer. One of her first responsibilities was TEX.

Cathy attended many of the TEX conferences in Europe (in Strasbourg, Karlsruhe and Cork), as well as TUG meetings in North America (in Montreal and Stanford), and some of the GUTenberg meetings in Paris. But it will be the TEX88 conference held at Exeter University with which she may be best remembered to many in the TEX world. Cathy was the local organiser for this meeting: the huge success of the meeting must be attributed in no small way to her vivacity and personal magnetism.

Cathy was an excellent teacher of TEX. She taught a number of TUG courses on Beginning and Intermediate TEX, and other TEX courses at Exeter University. At the University, she also developed courses which united her skill with TEX and her design abilities.

Besides her other activities in TEX and Electronic Publishing, she had been a member of the UK TEX Users’ Group committee since the foundation of the group, as well as a member of the committee of the BCS Electronic Publishing Specialist Group.

Malcolm Clark

Samuel B. Whidden

Sam Whidden, one of the founders of TUG, and its treasurer from 1980 through 1987, died unexpectedly on October 29, 1991. At the time of his death, he was a member of TUG’s Long Range Planning Committee.

Sam was a long-time employee of the American Mathematical Society. He established the Society’s information systems department in 1968, and was a champion of in-house composition of AMS publications, leading this function from the status of research project to a full production operation. He quickly saw the need for TEX users to band together for mutual support, and was instrumental in arranging for the first meeting at Stanford in 1980. Sam left regular employment at the Society in 1989, but continued as a consultant.

Barbara Beeton

TEX and TUG NEWS Table of Contents

Editorial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 First TUG Election by Mail . . . . . . . . . . . . . . . . . . . . . . . . 2 Announcement: Donald E. Knuth Scholarship for 1992 . . . . . . . . . 3 What is the L^{A}TEX3 Project?

Malcolm Clark . . . . . . . . . . . . . . . . . . . . . . . . . . 4 The L^{A}TEX3 Project Fund

Frank Mittelbach and Chris Rowley . . . . . . . . . . . . . . . 5 Scientific Publishing in Mexico

Kurt Bernardo Wolf . . . . . . . . . . . . . . . . . . . . . . . 7 Typographer’s Inn

Peter Flynn . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 New Publications . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 “Hey — it works!” . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

The Cauchy Integral Claudio Beccari . . . . . . . . . . . . . . . . 15 Dealing with Very Long Paragraphs Peter Schmitt . . . . . . . . 15

(L^{A})TEX News . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

IL^{A}TEX Joachim Schrod . . . . . . . . . . . . . . . . . . . . . . . . 16 News from Around . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

CyrTUG Irina Makhovaya . . . . . . . . . . . . . . . . . . . . . . 17 SibTUG Dr. Anatoly L. Urvantsev . . . . . . . . . . . . . . . . . 18

Reports on Meetings . . . . . . . . . . . . . . . . . . . . . . . . . . . . TUG’91 Christina Thiele . . . . . . . . . . . . . . . . . . . . . . 19 CsTUG Friedhelm Sowa . . . . . . . . . . . . . . . . . . . . . . . 20 EuroTEX91 Bernard Gaulle . . . . . . . . . . . . . . . . . . . . . 20 TECH’92 Denis Mégevand . . . . . . . . . . . . . . . . . . . . . . 21

TUG’92: TEX in Context

Mimi Lafrenz . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 TUG Board Activities . . . . . . . . . . . . . . . . . . . . . . . . . . .

Summary of Board Meeting . . . . . . . . . . . . . . . . . . . . . 23 Summary of Executive Committee Meeting . . . . . . . . . . . . 25

Upcoming Events . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 In Memoriam . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

Volume 1, No. 1, 1992

TEX in Context:

Resources, Support Tools, and Comparative Studies

July 26 to 30, 1992 Portland, Oregon 13th Annual TEX Users Group Meeting

The 1992 TEX Users Group Conference will focus on the TEX resources environment, on the tools used to support TEX, and on comparing and understanding the interaction of the various elements within the TEX world as well as with external products.

During four information-packed days, we’ll delve into the use of front ends for TEX, ways of including graphics in TEX documents, how to export TEX output to other graphics programs, comparisons of various personal computer implementations of TEX, network access and resources, educational issues, and translation between TEX and word-processors.

Mark your calendars and join us in Portland, the home of 20-pound salmon and 20-story buildings. Ride light rail trains over cobblestone streets, ski Mt. Hood and attend the symphony in the same day — even in July! Of special interest to TUG Meeting attendees may be the 11th Annual Mt. Hood Festival of Jazz to be held August 1st and 2nd in Gresham, Oregon, a suburb of Portland.

We’ll meet and stay at the Benson Hotel, a registered historic landmark recently restored to its former grandeur of the early 1900s. Special TUG rates: $89/night (available until 26 June only).

For a complete visitors’ guide, The Portland Book, call the Portland Visitors’ Center at (800) 345-3214.

Program Coordinator:  Program Committee
Mimi Jett ⋄ ETP Services Co. Helen Gibson ⋄ Wellcome Institute
For more information:  Doug Henderson ⋄ Blue Sky Research
tug@math.ams.com  Ron Whitney ⋄ TEX Users Group

Watch your mail and future issues of TUGboat and TEX and TUG NEWS for more details. Or contact the TUG Office. Meet you in Portland!